Use the definitions and theorems of this section to evaluate and simplify the following expression. Be sure to express answers with positive exponents. (b^3)^2
Answer by richard1234(5390)

Remember (b^3)^2 = (b^3)(b^3)

So (b^3)^2 = b^6